I am back to functions again.
The problem is stated as
y=25(-7x-4)^64 state the domain and range
I found the domain as "all real numbers" but can't figure out the range.
Who can help me please?

Answer by stanbon(75887)

The problem is stated as
y=25(-7x-4)^64 state the domain and range
I found the domain as "all real numbers" but can't figure out the range.
----------------------------------------
Comment: When you raise any Real Number to an even power the result
is always >= zero.
----------------------------------------
So, (-7x-4)^64 is never negative.
-7x-4 =0 when x = -4/7
---
The y-value when x = -4/7) is 25*0 = 0
----
So the lowest value y can have is zero.
----------------
Range is All Real Numbers >= 0
=====================================
Cheers,
Stan H.